# Long-term Frame-Event Visual Tracking: Benchmark Dataset and Baseline

### v1

@@ Abstract
Current event-/frame-event based trackers undergo evaluation on short-term tracking datasets, however, the tracking of real-world scenarios involves long-term tracking, and the performance of existing tracking algorithms in these scenarios remains unclear. In this paper, we first propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and has become the largest frame-event tracking dataset to date. We re-train and evaluate 15 baseline trackers on our dataset for future works to compare. More importantly, we find that the RGB frames and event streams are naturally incomplete due to the influence of challenging factors and spatially sparse event flow. In response to this, we propose a novel associative memory Transformer network as a unified backbone by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on both FELT and RGB-T tracking dataset LasHeR fully validated the effectiveness of our model. The dataset and source code can be found at https://github.com/Event-AHU/FELT_SOT_Benchmark.
@@ 1 Introduction
In this work, we also propose a novel frame-event tracking framework, termed AMTTrack. The key insight of our tracker is that the long-term tracking videos present the characteristics of incomplete information. For example, the RGB frames can be influenced by low illumination, motion blur, occlusion, etc, meanwhile, the event data is spatially sparse or with no target signal at all. This uncertain, incomplete, and not always complementary frame-event signal makes conventional fusion methods perform poorly. In this work, we build a unified backbone based on the modern Hopfield layer [19] which makes the ViT backbone have the ability of associative memory. Thus, the spatial-temporal information can be fully mined and learned. We integrate this strategy into the CEUTrack [22] and obtain higher performance on long-term frame-event tracking and RGB-T tracking (LasHeR dataset is evaluated). The overall framework is illustrated in Fig. 3.
@@ 5.1 Datasets and Evaluation Metric
For the evaluation of our tracker, we conduct our experiments on the long-term frame-event tracking dataset FELT. In addition, we also evaluate it on the RGB-Thermal tracking task using LasHeR [16] dataset to demonstrate the generalization of our tracking framework. LasHeR dataset is a large-scale benchmark dataset proposed for RGB-Thermal tracking, which contains 1,224 video sequences (pairs with 730K image pairs). It default divides the dataset into a training and a testing subset which contains 979 and 245 video sequence pairs.
@@ 6 Conclusion
In this work, we propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and is also the largest frame-event tracking dataset to date. To build a comprehensive benchmark, we also re-train and evaluate 15 baseline trackers on our newly proposed FELT SOT dataset, which will be a better platform for future works to compare. In addition, we find that the RGB frames and event streams are naturally incomplete data due to the influence of challenging factors and spatially sparse event data. To address this issue, in this work, we propose a novel associative memory Transformer network by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on multiple event-based tracking datasets fully validated the effectiveness of our model. In our future works, we will consider further improving the overall long-term tracking results by exploiting the dense temporal resolution of event data via asynchronous fusion modules.

### v2

@@ Abstract
Current event-/frame-event based trackers undergo evaluation on short-term tracking datasets, however, the tracking of real-world scenarios involves long-term tracking, and the performance of existing tracking algorithms in these scenarios remains unclear. In this paper, we first propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and has become the largest frame-event tracking dataset to date. We re-train and evaluate 15 baseline trackers on our dataset for future works to compare. More importantly, we find that the RGB frames and event streams are naturally incomplete due to the influence of challenging factors and spatially sparse event flow. In response to this, we propose a novel associative memory Transformer network as a unified backbone by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on RGB-Event (FELT), RGB-Thermal (RGBT234, LasHeR), and RGB-Depth (DepthTrack) datasets fully validated the effectiveness of our model. The dataset and source code can be found at https://github.com/Event-AHU/FELT_SOT_Benchmark.
@@ 1 Introduction
In this work, we also propose a novel frame-event tracking framework, termed AMTTrack. The key insight of our tracker is that the long-term tracking videos present the characteristics of incomplete information. For example, the RGB frames can be influenced by low illumination, motion blur, occlusion, etc, meanwhile, the event data is spatially sparse or with no target signal at all. This uncertain, incomplete, and not always complementary frame-event signal makes conventional fusion methods perform poorly. In this work, we build a unified backbone based on the modern Hopfield layer [20] which makes the ViT backbone have the ability of associative memory. Thus, the spatial-temporal information can be fully mined and learned. We integrate this strategy into the CEUTrack [23] and obtain higher performance on long-term frame-event tracking and RGB-Thermal (LasHeR, RGBT234 dataset), RGB-Depth (DepthTrack) tracking. The overall framework is illustrated in Fig. 3.
@@ 5.1 Datasets and Evaluation Metric
For the evaluation of our tracker, we conduct our experiments on the long-term frame-event tracking dataset FELT. In addition, we also evaluate it on the RGB-Thermal tracking task using LasHeR [17] and RGBT234 [15] dataset, and RGB-Depth tracking dataset (DepthTrack [30]) to demonstrate the generalization of our tracking framework. LasHeR dataset is a large-scale benchmark dataset proposed for RGB-Thermal tracking, which contains 1,224 video sequences (with 730K image pairs). It default divides the dataset into a training and a testing subset which contains 979 and 245 video sequence pairs. RGBT234 contains 234 pixel-level aligned RGB-Thermal video samples for tracking. The DepthTrack dataset are collected using Intel Realsense 415 event camera which output paired RGB and depth frames (with same resolution $640\times 360$, 30 FPS). It contains 150 and 50 training/validation sequences with 218,201 and 76,390 video frames, respectively.
@@ 5.3 Comparison on Public Benchmarks
Results on RGBT234 Dataset. As depicted in Table 5, we have compared our tracker with the latest state-of-the-art trackers on the extensive RGBT234 dataset, including APFNet, ProTrack, and others. For our performance evaluation, we utilize the MSR and MPR metrics. The comparison involved our proposed AMTTrck against recent RGB-T trackers, and as shown in Table 5, AMTTrck demonstrated remarkable results with a 58.0% MSR and 80.5% MPR. It secures the third position among the current state-of-the-art trackers. This experiment underscores the effectiveness and versatility of our method across diverse tracking tasks based on RGB+X.
Table 5: Tracking results on RGBT234 SOT dataset.
Trackers mfDiMP SGT DAFNet OSTrack FANet MaCNet
MPR 64.6 72.0 79.6 72.9 78.7 79.0
MSR 42.8 47.2 54.4 54.9 55.3 55.4
Trackers CAT JMMAC CMPP APFNet ProTrack Ours
MPR 80.4 79.0 82.3 82.7 79.5 80.5
MSR 56.1 57.3 57.5 57.9 59.9 58.0
Results on DepthTrack Dataset. As shown in Table 6, In our comparative analysis against leading trackers on the DepthTrack dataset, including DDiMP and DAL, the integration of Hopfield network layers has significantly enhanced our tracker’s performance. Specifically, our method has yielded impressive results with F-score, Recall, and Precision metrics reaching 48.6/49.7/47.6 respectively. Notably, our approach exhibits versatility by extending its effectiveness to RGB-D datasets as well. This experiment underscores the robustness and adaptability of our method across various RGB+X-based tracking tasks.
Table 6: Tracking results on DepthTrack SOT dataset.
Trackers SiamM_Ds Siam_LTD DAL LTMU_B ATCAIS DDiMP Ours
F-score 33.6 37.6 42.9 46.0 47.6 48.5 48.6
Recall 26.4 34.2 36.9 41.7 45.5 46.9 49.7
Precision 46.3 41.8 51.2 51.2 50.0 50.3 47.6
@@ 6 Conclusion
In this work, we propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and is also the largest frame-event tracking dataset to date. To build a comprehensive benchmark, we also re-train and evaluate 15 baseline trackers on our newly proposed FELT SOT dataset, which will be a better platform for future works to compare. In addition, we find that the RGB frames and event streams are naturally incomplete data due to the influence of challenging factors and spatially sparse event data. To address this issue, in this work, we propose a novel associative memory Transformer network by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on multiple multi-modal based tracking datasets (e.g., RGB-Event, RGB-Depth, RGB-Thermal) fully validated the effectiveness of our model. In our future works, we will consider further improving the overall long-term tracking results by exploiting the dense temporal resolution of event data via asynchronous fusion modules.